OpenStudy (icalibear):

Simplify to just one term (1−5/6x)÷1x−5/6x2

OpenStudy (icalibear):

\[(1- \frac{ 5 }{ 6x })\div \frac{ 1 }{ x}-\frac{ 5 }{ 6x^2 }\]

OpenStudy (anonymous):

Try a common denominator of 6x for the terms in parentheses then subtract

OpenStudy (icalibear):

I think I'm supposed to flip the second one and multiply across?

OpenStudy (icalibear):

I also feel like some things would cancel?

OpenStudy (anonymous):

when you divide, yes. but first you have to handle the subtraction in parentheses

OpenStudy (icalibear):

Actually there are no parentheses in the question I just put them there to separate the two

OpenStudy (icalibear):

ugh

OpenStudy (anonymous):

or is the problem supposed to be